OpenStudy (anonymous):

Which of the following is a solution for the system y = -3x + 5 and y = 2x + 5?

OpenStudy (anonymous):

x=0 ?

OpenStudy (radar):

y=y -3x+5=2x+5 -5x=0 x=0/-5=0 y=5